# MotionCLR: Motion Generation and Training-free Editing via Understanding Attention Mechanisms

### v1

@@ 5.3 Ablation Study
We provide some ablation studies on some technical designs. (1) The setting w/o separate word modeling shows poorer qualitative results with the w/ separate word setting. The separate word-level cross-attention correspondence benefits better text-to-motion controlling, which is critical for motion fine-grained generation. (2) The setting of injecting text tokens before motion tokens performs worse than the MotionCLR. This validates the effectiveness of introducing the cross-attention for cross-modal correspondence. The ablation studies additionally verify the basic motivation of modeling word-level correspondence in MotionCLR.
5.4 Action Counting from Attention Map
@@
6 Failure Cases Analysis and Correction
[…]
@@ 6 Failure Cases Analysis and Correction
There are few generative methods that can escape the curse of hallucination. In this section, we will discuss some failure cases of our method and analyze how we can refine these results. The hallucination of counting is a notoriously tricky problem for generative models, attracting significant attention in the community and lacking a unified technical solution. Considering that this problem cannot be thoroughly resolved, we try to partially reveal this issue by additionally providing temporal grounds. For example, if the counting number of an action is not aligned with the textual prompt, we can correct this by specifying the temporal grounds of actions. Technically, the temporal mask can be treated as a sequence of weights to perform the motion emphasizing and de-emphasizing. Therefore, grounded motion generation can be easily achieved by adjusting the weights of words.
@@
Specifically, we show some failure cases of our method. As shown in Fig. 13, the generated result of “a person jumps four times” fails to show four times of jumping actions, but seven times. To meet the requirement of counting numbers in the textual prompts, we additionally input a temporal mask, including four jumping timesteps, to provide temporal grounds. From the root height visualization and the motion visualization, the times of the jumping action have been successfully corrected from seven to four. Therefore, our method is promising for grounded motion generation to reveal the hallucination of deep models.
Moreover, other editing fashions are also potential ways to correct hallucinations of generated results. For example, the motion sequence shifting and in-place motion replacement functions can be used for correcting sequential errors and semantic misalignments, respectively.
7 Conclusion and Future Work
Conclusion. In this work, we propose a diffusion-based motion generation model, MotionCLR. With this model, we carefully clarify the self-attention and the cross-attention mechanisms in the MotionCLR. Based on both theoretical and empirical analysis of the attention mechanisms in MotionCLR, we developed versatile motion editing methods. Additionally, we not only verify the action counting ability of attention maps, but also show the potential of motion corrections. We build a user interface in Appendix B to demonstrate how our method can support interactive editing.
Limitation and future work. As shown in Sec. 6, our model can also not escape the hallucination curse of generative models. Therefore, we leave the grounded motion generation as future work. As discussed in Lu et al. (2024), the CLIP model used in MotionCLR is still a bit unsatisfactory. Therefore, we will provide token-level text-motion alignment encoders to provide textual conditions. Besides, our future work will also focus on scaling the MotionCLR to generate more vivid motions.
Border impact statement. The development of MotionCLR, a diffusion-based motion generation model, has the potential to impact various fields of motion synthesis and editing significantly. However, the complexity of the MotionCLR and its performance limitations under certain conditions may lead to errors or inaccuracies in practical applications. This could result in negative consequences in critical fields such as humanoid simulation and autonomous driving. Therefore, it is necessary to further optimize the model and carefully assess its reliability before widespread deployment.
Acknowledgement
The author team would like to acknowledge Dr. Jingbo Wang from Shanghai AI Laboratory and Dr. Xingyu Chen from Peking University for his constructive suggestions and discussions on downstream applications. We also would like to acknowledge Mr. Hongyang Li and Mr. Zhenhua Yang from SCUT for their detailed discussion on some technical details and writing. Mr. Bohong Chen from ZJU also provided us with insightful feedback on the evaluation and the presentations. We convey our thanks to all of them.
Contribution Statement
Ling-Hao Chen contributes to the key design of MotionCLR, code implementation, and visualization. Wenxun Dai and Shunlin Lu contribute to motion kinematics and visualization to accelerate the progress of this project. Xuan Ju contributes to the methods and technical implementations of some motion editing applications. Lei Zhang is the advisor throughout the project. All authors commit to the writing of this paper.
@@ Acknowledgement
Appendix A Supplemental Experiments
A.1 What is the Self-attention Map like in Motion (De-)emphasizing?
This experiment is an extension of the experiment shown inm Fig. 5.
[…]
Compared to Fig. 14a, Fig. 14b shows two fewer trajectories. This reduction is due to the de-emphasizing effect, where the character’s second jump was not fully executed, resulting in just an arm motion (Fig. 5(b)). Consequently, the two actions became distinguishable, leading to fewer detected two trajectories. In Fig. 14c, the second jumping has been completely erased, resulting in only one trajectory, further demonstrating how de-emphasizing significantly affects motion execution.
[…]
A.2 What is the Difference between Motion (De-)emphasizing in MotionCLR and Adjusting Classifier-free Guidance Weights?
In this part, we would like to discuss the difference between reweighting the cross-attention map and adjusting classifier-free guidance weights.
As shown in Tab. 4, we ablate how different $w$s affect the results. The results suggest that adjustment of $w$ impacts the quality of generated results, making $w=2.5$ an effective choice. When $w$ increases, the text-motion alignment increases consistently, and the generation quality (FID) requires a trade-off.
$w$ 1 1.5 2 2.5 3 3.5
FID 0.801 0.408 0.318 0.217 0.317 0.396
TMR-sim. 51.987 52.351 53.512 53.956 54.300 54.529
Table 4: Differnent editing results when changing $w$s. In MotionCLR, $w=2.5$ is the default design choice for the denoising sampling.
However, as the classifier-free guidance mainly works for the semantic alignment between text and motion, it cannot control the weight of each word. We take the “a man jumps.” as an example for a fair comparison, which is the case used in the main text [^1] Suggested to refer to Fig. 5 for comparison.. As shown in Fig. 15, the generated motions with different $w$ values illustrate that $w$ cannot influences both the height and frequency of the jump. Nevertheless, the classifier-free guidance is limited in its ability to control more detailed aspects, such as the exact height and number of actions. Therefore, while $w$ improves text-motion alignment, it cannot achieve fine-grained adjustments.
A.3 More Experimental Results of In-place Motion Replacement
Semantic similarity of edited motions. In the in-place motion replacement application, we measure the editing quality and the text-motion similarity. To verify this, we construct a set of prompts, tagged with the edited words. In Tab. 5, we compare our method (ours replaced) with the unedited ones (unreplaced) and the generated motions directly changing prompts (pseudo-GT). As can be seen in Tab. 5, the motions of the three groups have similar qualities. Besides, the edited motion is similar to the pseudo-GT group, indicating the good semantic alignment of the edited results.
FID $\downarrow$ TMR-sim.$\rightarrow$
direct (pseudo GT) 0.315 0.543
unreplaced 0.325 0.567
unreplaced (unpaired T-M) 0.925 0.490
ours replaced 0.330 0.535
Table 5: Comparison between the generation result with directly changing prompts and the in-place replacement in MotionCLR. The semantics of editing results are similar to the motion directly generated by the changed prompt. The setting difference between “unreplaced” with “unreplaced (unpaired T-M)” is that the latter texts are edited sentences.
Ablation study of different attention layers. To further explore the impact of attention manipulation in in-place motion replacement, we conduct an ablation study by varying the layers in MotionCLR for manipulation, shown in Tab. 6. The table lists the results for different ranges of manipulated attention layers. It can be observed that manipulating different attention layers influences the editing quality and the semantic similarity (TMR-sim.). In particular, manipulating the layers from 1 to 18 achieves the best semantic consistency, demonstrating the effectiveness of editing across multiple attention layers for maintaining semantic alignment in the edited motion. The less effectiveness of manipulating middle layers is mainly due to the fuzzy semantics present in the middle layers of the U-Net. As these layers capture more abstract with reduced temporal resolution, the precise details and localized information become less distinct. Consequently, manipulating these layers has a limited impact on the final output, as they contribute less directly to the fine-grained details for the task.
[…]
8 11 0.339 0.472
5 14 0.325 0.498
1 18 0.330 0.535
Table 6: The ablation study of manipulating different attention layers. The “begin” and “end” represent the beginning and the final layer for manipulation.
A.4 Comparison with Manipulation Noisy Motions in the Diffusion Process
As the diffusion denoising process can manipulate the motion directly in the denoising process, this is a baseline for comparison with our motion shifting and example-based motion generation applications. Here, for convenience, we only take the example-based motion generation application as an example for discussion. In this section, we conduct a comparison between our proposed editing method and diffusion manipulation in the motion space, focusing on the FID and diversity metrics. The 200 samples used in this experiment were constructed by researchers. As depicted in Tab. 7, the “Diff. manipulation” serves for our comparison. Our method achieves an FID value of 0.427, indicating a relatively high generation quality, while the “Diff. manipulation” achieves a higher FID of 0.718, demonstrating worse fidelity. Conversely, in terms of diversity, the “MotionCLR manipulation” exhibits a higher diversity (Div.) score of 2.567 compared to the 1.502 of the “Diff. manipulation.” These results verify our method is better than manipulating noisy motions in the denoising process. The main reason for the better quality and diversity mainly relies on the many times of manipulation of self-attention, but not the motion. Directly manipulating the motion results in some jitters, making more effort for models to smooth. Besides, the shuffling times of manipulating the self-attention maps are higher than the baseline, contributing to the better diversity of our method.
FID $\downarrow$ Div. $\uparrow$
Diff. manipulation 0.718 1.502
MotionCLR manipulation 0.427 2.567
Table 7: Comparison on FID and diversity values with manipulating self-attention in the motion space of the denoising process.
A.5 Motion Generation Result Visualization
@@
Figure 16: Human motion generation result of MotionCLR.
A.6 More Visualization Results of Motion Sequence Shifting
@@
A.7 More Visualization Results on Exampel-based Motion Generation
We provide some visualization results to further illustrate the effectiveness of our approach in generating diverse motions that adhere closely to the given prompts. In Fig. 18, the example motion of “a person kicking their feet” is taken as the reference, and multiple diverse kick motions are generated. These generated motions not only exhibit variety but also maintain key characteristics of the original example. Similarly, in Fig. 19, the example motion of “a person walking in a semi-circular shape while swinging arms slightly” demonstrates the capability to generate diverse walking motions that maintain the distinct features of the source motion. The generated trajectories, in Fig. 18(b) and Fig. 19(b), show that diverse motions follow different paths while retaining similarities with the original one, confirming the effectiveness of our method.
@@ A.7 More Visualization Results on Exampel-based Motion Generation
(a) The example motion (blue) and the generated diverse motion (red).
Figure 19: Diverse generated results of blue example generated by the prompt “person walks in a semi circular shape while swinging arms slightly.”. The example-based generated walking motions are diverse and similar to the source walking example.
A.8 Detailed Visualization Results of Grounded Motion Genration
(a) The root height comparison. The red area denotes the timesteps to execute actions.
[…]
As depicted in Fig. 20, we provide a detailed comparison between the motion generation results with and without grounded settings. While the main text (Sec. 6) has already discussed the general differences between these settings, here in the appendix, we further extract and visualize the root height trajectory separately for a clearer and more detailed comparison. This approach helps in highlighting the effectiveness of our method in addressing motion hallucination issues and ensuring that the generated movements closely align with the given textual prompts.
In Fig. 20(a), the root height comparison distinctly shows the difference between the edited and vanilla results. The red-shaded regions indicate the time steps where the specified actions (“jumps four times”) should occur. Without grounded motion generation, the vanilla result tends to generate more than the required number of jumps, resulting in motion hallucination. However, with the incorporation of temporal grounding, our edited result accurately performs the action four times, aligning with the textual prompt. Fig. 20(b) further visualizes the motion sequences. It is evident that the temporal grounding guides the motion generation process, ensuring consistency with the input prompt. The edited result follows the correct sequence of actions, demonstrating the advantage of using grounded motion settings to avoid common hallucinations in generative models.
Overall, these detailed visualization results confirm the importance of incorporating temporal grounding into motion generation tasks, as it helps mitigate hallucinations in generative models, ensuring the generated motions are more faithfully aligned with the intended textual descriptions.
Appendix B User Interface for Interactive Motion Generation and Editing
To have a better understanding of our task, we build a user interface with Gradio [Abid et al., 2019]. We introduce the demo as follows.
In Fig. 21, we illustrate the steps involved in generating and visualizing motions using the interactive interface. Fig. 21(a) displays the initial step where the user provides input text such as “a man jumps” and adjusts motion parameters. Once the settings are finalized, the system begins processing the motion based on these inputs, as seen in the left panel. Fig. 21(b) showcases the generated motion based on the user’s input. The interface provides a rendered output of the skeleton performing the described motion. This presentation allows users to easily correlate the input parameters with the resulting animation. The generated motion can further be edited by adjusting parameters such as the length of the motion, emphasizing or de-emphasizing certain actions, or replacing actions altogether, depending on user requirements. This process demonstrates how the interface facilitates a workflow from input to real-time motion visualization.
@@ C.1 Mathematical Visualization of Self-attention Mechanism
In the main text (Eq. 2), we introduced the self-attention mechanism of MotionCLR, which utilizes different transformations of motion as inputs. The motion embeddings serve as both the query ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}Q}$), key ($\mathbf{\color[rgb]{0.9765625,0.828125,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9765625,0.828125,0.703125}K}$), and value ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$), capturing the internal relationships within the sequence of motion frames.

### v2

@@ 5.6 Ablation Study
Table 8: The ablation study of manipulating different attention layers on the HVerb-wild test set. The “begin” and “end” represent the beginning and the final layer/step for manipulation. The bottom row denotes our design choice for motion editing.
[…]
begin end begin end
1 9 8 11 0.339 63.5 62.5
1 9 5 14 0.335 65.7 59.2
3 7 1 18 0.399 64.7 59.3
4 6 1 18 0.455 62.8 59.0
1 9 1 18 0.330 66.0 58.1
In this section, we provide ablation studies for the generation function and the editing functions, respectively.
Ablation study of the generation ability in MotionCLR. We provide some ablation studies on some technical designs in Tab. 7. (1) The setting w/o separate word modeling shows poorer qualitative results with the w/ separate word setting. The separate word-level cross-attention correspondence benefits better text-to-motion controlling, which is critical for motion fine-grained generation. (2) The setting of injecting text tokens before motion tokens performs worse than the MotionCLR. This validates the effectiveness of introducing the cross-attention for cross-modal correspondence. The ablation studies additionally verify the basic motivation of modeling word-level correspondence in MotionCLR.
Ablation study of editing different attention layers. To further explore the impact of attention manipulation, we conduct an ablation study by varying the layers in MotionCLR for manipulation, shown in Tab. 8. Without losing generalization, we test this on the in-place motion replacement application. The table lists the results for different ranges of manipulated attention layers. It can be observed that manipulating different attention layers influences the editing quality and the semantic similarity (TMR-sim.). In particular, manipulating the layers from 1 to 18 achieves the best semantic consistency, demonstrating the effectiveness of editing across multiple attention layers for maintaining semantic alignment in the edited motion. The less effectiveness of manipulating middle layers is mainly due to the fuzzy semantics present in the middle layers of the U-Net. As these layers capture more abstract with reduced temporal resolution, the precise details and localized information become less distinct. Consequently, manipulating these layers has a limited impact on the final output, as they contribute less directly to the fine-grained details of the task.
[…]
5.7 User Study
Table 9: User study results. The study evaluates the motion (de-)emphasis and in-place replacement quality, respectively. The results show that our method outperforms other baselines across fidelity, user requirement satisfaction, and content similarity with the unedited motion.
(de-)emphasis fidelity $\uparrow$ satisfactory $\uparrow$ similarity $\uparrow$
MotionFix 3.51 3.80 4.07
Ours 4.02 3.95 4.25
replacement fidelity $\uparrow$ satisfactory $\uparrow$ similarity $\uparrow$
editing text only 3.04 3.55 3.01
MotionFix 3.10 2.78 3.22
Ours 4.05 4.26 3.99
To additionally verify the effectiveness of the semantic editing result, we also compare our method with baselines via human evaluation. Here, for the semantic motion editing method, we mainly focus on the motion (de-)emphasis and the im-place motion replacement task, both of which are based on editing cross-attention.
[…]
5.8 Potential Action Counting Ability from Attention Map
@@
6 Case Study: Failure Cases Analysis and Correction
@@
There are few generative methods that can escape the curse of hallucination. In this section, we will discuss some failure cases of our method and analyze how we can refine these results. The hallucination of counting is a notoriously tricky problem for generative models, attracting significant attention in the community and lacking a unified technical solution. Considering that this problem cannot be thoroughly resolved, we try to partially reveal this issue by additionally providing temporal grounds. For example, if the counting number of an action is not aligned with the textual prompt, we can correct this by specifying the temporal grounds of actions. Technically, the temporal mask can be treated as a sequence of weights to perform the motion emphasis and de-emphasis. Therefore, grounded motion generation can be easily achieved by adjusting the weights of words.
@@
Specifically, we show some failure cases of our method. As shown in Fig. 14, the generated result of “a person jumps four times” fails to show four times of jumping actions, but seven times. To meet the requirement of counting numbers in the textual prompts, we additionally input a temporal mask, including four jumping timesteps, to provide temporal grounds. Technically, for the desired four jumping areas, we set $\alpha=0.15$ as emphasis. For the reserving area, we set $\alpha=-0.1$ as de-emphasis. From the root height visualization and the motion visualization, the times of the jumping action have been successfully corrected from seven to four. Therefore, our method is promising for grounded motion generation to reveal the hallucination of deep models. Moreover, other editing fashions are also potential ways to correct hallucinations of generated results. For example, the motion sequence shifting and in-place motion replacement functions can be used for correcting sequential errors and semantic misalignments, respectively.
(a) Web interface for MotionCLR.
(b) Blender add-on for MotionCLR.
Figure 16: Web interface and the Blender add-on of MotionCLR, supporting proposed applications. In these two interfaces, we take the motion (de-)emphasis application as an example. Both interfaces support adjusting the weight of parameters to refine the result, until reaching satisfaction.
7 User Interface
MotionCLR is a simple yet effective framework that can be integrated into industrial tools. We develop a web interface of the MotionCLR model for interactive motion generation and editing. The web interface provides a quick view for new customers without installing any dependencies. For experienced users, like animators, we also support a Blender add-on of MotionCLR similar to the web demo, supporting the real creation loop of motion synthesis. Two interfaces are shown in Fig. 16(a) and Fig. 16(b), respectively. In Fig. 16, both the web interface and add-on support all applications of MotionCLR. Besides, users are allowed to interactively adjust the weight of parameters to refine the result, until reaching their satisfaction.
8 Discussion and Conclusion
In this work, we carefully clarify what roles cross-attention plays in motion generation, enhancing the explainability of attention mechanisms in text-to-motion generation. Except for the comparable generative performance with SOTA methods, our proposed MotionCLR model supports diverse interactive motion generation and editing in one system for the first time. Importantly, in the editing process, MotionCLR supports editing a motion without any training, which even outperforms some methods requiring specific training.
To have a thorough exploration of the proposed method, we construct an evaluation set to validate the theoretical analysis of the attention mechanism. Based on these preliminary experiments and new understanding, we evaluate our applications with specific baselines. Our method even outperforms some specifically designed methods in some scenarios. Besides, we also explore the boundaries of our method in action counting and grounded motion generation. Considering real-world applications in the animation creation, we develop a web interface and a Blender add-on for users.
This work explores the new interaction fashion of motion synthesis. We are still facing some limitations, driving for developing future methods. 1) Different interaction fashions. This work introduces a new interaction fashion for motion generation, relying on multiple interactions with the machine. However, the language-based chatting interaction is also useful to users, which is leaving as our future work. Besides, MotionCLR can also serve as a data generator to synthesize a dataset for such applications. 2) Robustness of the method. As shown in Sec. 6, our model can also not escape the hallucination curse of generative models. Although the proposed grounded motion generation method can relieve the phenomena, it is still worth enhancing the capability of the base model. To this end, the grounded synthesis method could provide additional supervision in this process. We will reach this in the future. Besides, our method will also benefit from a larger dataset with good quality, which is already in the collecting process before submission.
@@ 8 Discussion and Conclusion
Appendix A Supplemental Experiments
A.1 What is the Self-attention Map like in Motion (De-)emphasis?
@@
A.2 Motion Generation Result Visualization
@@
A.3 More Visualization Results of Motion Sequence Shifting
@@
A.4 More Visualization Results on Exampel-based Motion Generation
We provide some visualization results to further illustrate the effectiveness of our approach in generating diverse motions that adhere closely to the given prompts. In Fig. 21, the example motion of “a person kicking their feet” is taken as the reference, and multiple diverse kick motions are generated. These generated motions not only exhibit variety but also maintain key characteristics of the original example. Similarly, in Fig. 21, the example motion of “a person walking in a semi-circular shape while swinging arms slightly” demonstrates the capability to generate diverse walking motions that maintain the distinct features of the source motion. The generated trajectories, as visualized in Fig. 20(b) and Fig. 20(d), show that the diverse motions follow different paths while retaining similarities with the original motion, confirming the effectiveness of our method.
A.5 Inversion v.s. Generation
Our method can directly adopt DDIM inversion to edit the existing GT motions. Here, we compare the generation results with DDIM output. As shown in Tab. 10, the inversion result of the generated motion is similar to the generated motion, supporting good performance of editing GT motions.
Table 10: Comparison between the inversed and generated motions.
[…]
Top 1 Top 2 Top 3
generated 0.544 0.732 0.831 0.269
generated + inversion 0.535 0.724 0.818 0.299
Figure 18: More visualization of human motion generation result by MotionCLR.
A.6 Results on Other Datasets
We further evaluate our method on the KIT-ML dataset Plappert et al. [2016]. As shown in Sec. A.5, MotionCLR outperforms the state-of-the-art MoMask in key metrics. These results demonstrate the robustness and generalizability of MotionCLR, with consistent performance trends across datasets of different sizes.
Table 11: Comparison of results on KIT-ML. MotionCLR demonstrates superior performance on key metrics.
Method Top 1 Top 2 Top 3 FID MM-Dist Multi-Modality
MoMask (SOTA) 0.433 0.656 0.781 0.204 2.779 1.131
MotionCLR 0.438 0.658 0.783 0.275 2.773 1.213
@@ A.6 Results on Other Datasets
(c) The example motion (blue) and the generated diverse motion (red).
(d) The trajectory visualizations of the example motion and diverse motions.
Figure 21: Diverse generated results of blue example generated by the prompt “person walks in a semi-circular shape while swinging arms slightly.”. The example-based generated walking motions are diverse and similar to the source walking example.
Appendix B Web User Interface for Interactive Motion Generation and Editing
To have a better understanding of our task, we build a user interface with Gradio [Abid et al., 2019]. We introduce the demo as follows. In Fig. 31, we illustrate the steps involved in generating and visualizing motions using the interactive interface. Fig. 31(a) displays the initial step where the user provides input text such as “a man jumps” and adjusts motion parameters. Once the settings are finalized, the system begins processing the motion based on these inputs, as seen in the left panel. Fig. 31(b) showcases the generated motion based on the user’s input. The interface provides a rendered output of the skeleton performing the described motion. This presentation allows users to easily correlate the input parameters with the resulting animation. The generated motion can further be edited by adjusting parameters such as the length of the motion, emphasizing or de-emphasizing certain actions, or replacing actions altogether, depending on user requirements. This process demonstrates how the interface facilitates a workflow from input to motion visualization.
@@ C.1 Mathematical Visualization of Self-attention Mechanism
In the main text (Eq. 2), we introduced the self-attention mechanism of MotionCLR, which utilizes different transformations of motion as inputs. The motion embeddings serve as both the query ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}Q}$), key ($\mathbf{\color[rgb]{0.9765625,0.828125,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9765625,0.828125,0.703125}K}$), and value ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$), capturing the internal relationships within the sequence of motion frames. Fig. 25 provides a detailed mathematical visualization of this process: (1) Similarity Calculation. In the first step, the similarity between the motion embeddings at different frames is computed using the dot product, represented by $\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}=\mathbf{{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}Q}{\color[rgb]{0.9765625,0.828125,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9765625,0.828125,0.703125}K}}^{\top}$. This measurement reflects the internal relationship/similarity between different motion frames within the sequence. Fig. 24(a) illustrates how the $\mathtt{softmax}(\cdot)$ operation is applied to the similarity matrix to determine which motion feature should be selected at a given frame $f$. (2) Feature Updating. Next, the similarity scores are used to weight the motion embeddings ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$) and generate updated features $\mathbf{\color[rgb]{0.92578125,0.71484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.71484375,0.515625}X^{\prime}}$, as shown by the equation $\mathbf{\color[rgb]{0.92578125,0.71484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.71484375,0.515625}X^{\prime}}=\mathtt{softmax}(\mathbf{{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}Q}{\color[rgb]{0.9765625,0.828125,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9765625,0.828125,0.703125}K}}^{\top}/\sqrt{d})\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$. Here, the similarity matrix applies its selection of values ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$) to update the motion features. This process allows the self-attention mechanism to dynamically adjust the representation of each motion frame based on its relevance to other frames in the sequence.
C.2 Mathematical Visualization of Cross-attention Mechanism
(1) Similarity Calculation. In the first step, the similarity between the motion embeddings ($\mathbf{{\color[rgb]{1,0.6015625,0.24609375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.24609375}Q}}$) with $F$ frames and the text embeddings ($\mathbf{{\color[rgb]{0.94921875,0.62109375,0.66015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94921875,0.62109375,0.66015625}K}}$) with $N$ words is computed through the dot product, represented by $\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}=\mathbf{{\color[rgb]{1,0.6015625,0.24609375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.24609375}Q}{\color[rgb]{0.94921875,0.62109375,0.66015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94921875,0.62109375,0.66015625}K}}^{\top}$. This similarity measurement reflects the relationship between motion frames and words. Fig. 24(c) shows how the $\mathtt{softmax}(\cdot)$ operation is applied to the similarity matrix to determine which word should be activated at a given frame $f$.
As discussed in the main text (see Sec. 1), despite the progresses in human motion generation [Zhang et al., 2024c, Cai et al., 2024, Zhang et al., 2024b, Guo et al., 2024c, Raab et al., 2024b, Kapon et al., 2024, Cohan et al., 2024, Fan et al., 2024, Xu et al., 2023a, b, Yao et al., 2022, Feng et al., 2023, Ao et al., 2023, Yao et al., 2024, Zhang et al., 2024d, Liu et al., 2010, Aberman et al., 2020a, Karunratanakul et al., 2024, Li et al., 2024, 2023a, Gong et al., 2023, Zhou and Wang, 2023, Zhong et al., 2023, Athanasiou et al., 2023, Zhong et al., 2024, Guo et al., 2024b, Zhang et al., 2024b, Zhao et al., 2023, Zhang et al., 2022, 2020, Diomataris et al., 2024, Pinyoanuntapong et al., 2024, Diller and Dai, 2024, Peng et al., 2023, Hou et al., 2023, Liu et al., 2023, Cong et al., 2024, Jiang et al., 2022, Kulkarni et al., 2024, Tessler et al., 2024, Liang et al., 2024, Ghosh et al., 2023, Wu et al., 2024], there still lacks a explicit modeling of word-level cross-modal correspondence in previous work. To clarify this, our method models a fine-grained word-level cross-modal correspondence.
As illustrated in Fig. 26, the major distinction between our proposed method and previous diffusion-based motion generation models lies in the explicit modeling of word-level cross-modal correspondence. In the MDM-like fashion Tevet et al. [2022b] (see Fig. 26a), previous methods usually utilize a denoising transformer encoder that treats the entire text as a single embedding, mixing it with the motion sequence. This approach lacks the ability to capture the nuanced relationship between individual words and corresponding motion elements, resulting in an over-compressed representation. Although we witness that Zhang et al. [2024a] also introduces cross-attention in the motion generation process, it still faces two problems in restricting the fine-grained motion editing applications. First of all, the text embeddings are mixed with frame embeddings of diffusion, resulting in a loss of detailed semantic control. Our approach disentangles the diffusion timestep injection process in the convolution module to resolve this issue. Besides, the linear cross-attention in MotionDiffuse is different from the computation process of cross-attention, resulting in a lack of explanation of the word-level correspondence. The auto-regressive (AR) fashion [Zhang et al., 2023a] (Fig. 26b) adopts a simple concatenation of text and motion, where an AR transformer processes them together. However, this fashion also fails to explicitly establish a fine-grained correspondence between text and motion, as the AR transformer merely regards the text and motion embeddings as one unified sequence.